# Red Queen: Safeguarding Large Language Models against Concealed Multi-Turn Jailbreaking

### v1

@@ 7 Conclusions
We introduce Red Queen Attack, the first jailbreak method that constructs multi-turn scenarios to conceal harmful intent by claiming to prevent others from conducting malicious behavior. We develop 40 scenarios based on occupation and relationship with different lengths, combined with 14 categories of harmful actions, resulting in a dataset of 56k high-quality multi-turn attack examples. Our evaluation across ten models from five major model families confirms the effectiveness of the Red Queen Attack. Further analysis reveals that both concealment and multi-turn structures enhance the attack’s effectiveness, with larger models being particularly vulnerable. We also introduce a mitigation strategy called RED QUEEN GUARD, which aligns LLMs to effectively counter adversarial attacks. This approach reduces the attack success rate to below 1% while maintaining the model’s performance across standard benchmarks.
@@
Additionally, we focus on constructing multi-turn scenarios extensively in the text domain without exploring their effectiveness in multi-modality contexts, such as visual or speech. The introduction of a new modality can introduce orthogonal vulnerabilities, such as perception (Zhang et al., 2024) or modality alignment (Li et al., 2024). Therefore, our conclusions are primarily applicable to text-based jailbreak research.
Ethical Considerations
This study is centered on exploring the potential security vulnerabilities in large language models (LLMs) and mitigating them. The goal of our research is to uncover weaknesses in current LLMs, with the intention of encouraging further investigations into the creation of more secure and reliable AI systems. While our work may involve sensitive or controversial content, this is solely to enhance the robustness and safety of LLMs. The research findings are solely intended for academic purposes and must not be misused.
@@ Appendix A Harmful Plans Output
We provide the one result of Red Queen attack on ChatGPT web interface [^5] https://chat.openai.com/ in Figure 6.
Content Warning: This section contains examples of harmful language and plans.
@@
[…]
@@
[…]
We show Red Queen Attack success rates on different scenarios and harmful action categories in Tables 9, 10, 11 and 12. We highlight the best ASR in bold and the second in underline.
@@ Appendix B Red Queen result on different scenario and harmful action categories
[…]
@@
[…]
@@
[…]
Figure 7: Heatmap for GPT-4o result in five turn Red Queen Attack
Figure 8: Heatmap for Llama3-70b result in three turn Red Queen Attack
Figure 9: Heatmap for Mixtral-22b result in five turn Red Queen Attack
Appendix C Data Sample
Table 13 shows harmful action samples for each category. Tables 14, 15 and 16 shows the average token length of different turns.
Table 13: Harmful action categories and examples.
@@ Appendix C Data Sample
Table 14: Average token length of three-turn scenario.
@@
Table 15: Average token length of four-turn scenario.
@@
Table 16: Average token length of five-turn scenario.
@@
Appendix D Judgment Function Comparison
D.1 Details of Judgment Methods
We demonstrate details of the judgment method listed in Table 2 in the following:
Table 17: GCG Keywords
@@ D.1 Details of Judgment Methods
Bert-based Classifier (Huang et al., 2024b) is a fine-tuned classifier trained on using HH-RLHF dataset (Ganguli et al., 2022). The classifier estimates a risk score between 0 to 1 for each response. The attacks will be classified as successful if the risk exceeds 0.5.
System Judging Prompt is designed to label responses as harmful or not. It works well with GPT-4 and Llama3. This prompt evaluates whether the response includes a detailed plan or hints for creating a plan. We recommend testing this prompt first when adapting it to other tasks. The prompt is shown in Table 14, we evaluate the attack is successful if response contains “[Yes]”.
@@
D.2 Failure Example of Previous Evaluators
Table 20: Failure example of GCG. The red word is involved in the keyword list in Table 17
@@ D.2 Failure Example of Previous Evaluators
Table 23: Failure example of Llama Guard Judgment Method.
@@
Appendix E DPO Alignment Experiment
Appendix F Prompt Template
We provide actual Red Queen Attack of different scenarios in this section. Due to ethical considerations, we only provide all scenarios (LABEL:tab:attack_scenario) in three turns and police scenarios in all turns. The full attack is available on https://github.com/kriti-hippo/red_queen.

### v2

@@ 7 Conclusions
We introduce Red Queen Attack, the first jailbreak method that constructs multi-turn scenarios to conceal harmful intent by claiming to prevent others from conducting malicious behavior. We develop 40 scenarios based on occupation and relationship with different lengths, combined with 14 categories of harmful actions, resulting in a dataset of 56k high-quality multi-turn attack examples. Our evaluation across ten models from four major model families confirms the effectiveness of the Red Queen Attack. Further analysis reveals that both concealment and multi-turn structures enhance the attack’s effectiveness, with larger models being particularly vulnerable. Additionally, we propose Red Queen Guard, a mitigation strategy that reduces the attack success rate to below 1% while preserving model performance on standard benchmarks. Future work will focus on expanding the dataset’s diversity and exploring mitigation strategies through Reinforcement Learning from Human Feedback or prompt-based defences with intention detection.
@@
Additionally, we focus on constructing multi-turn scenarios extensively in the text domain without exploring their effectiveness in multi-modality contexts, such as visual or audio. The introduction of a new modality can introduce orthogonal vulnerabilities, such as perception (Zhang et al., 2024; Jiang et al., 2024c) or modality alignment (Li et al., 2024; Yang et al., 2025a), which further complicate the model’s ability to remain both helpful and harmless, particularly in reasoning-intensive scenarios (Jiang et al., 2023a; Yang et al., 2025b) where subtle intent or contextual cues play a crucial role. Therefore, our conclusions are primarily applicable to text-based jailbreak research.
Ethical Statement
This study is centered on exploring the potential security vulnerabilities in large language models (LLMs) and mitigating them. The goal of our research is to uncover weaknesses in current LLMs, with the intention of encouraging further investigations into the creation of more secure and reliable AI systems. While our work may use and involve sensitive or controversial content, this is solely to enhance the robustness and safety of LLMs. The research findings are solely intended for academic purposes and must not be misused.
@@ Appendix A Harmful Plans Output
We provide the one result of Red Queen Attack on ChatGPT web interface [^2] https://chat.openai.com/ in Figure 6. We provide two failure cases of Red Queen Attack in Figure 7.
@@
Figure 7: Example of failure cases
We provide more examples of harmful plan models generated in LABEL:tab:model_output and random sentences generated from small models, Mixtral-8$\times$7B and Qwen-7B, in Tables 8 and 9.
We further break down the results into different scenarios and harmful action categories in Figure 8. Red Queen Attack can have different ASR based on the scenario and action categories. Among occupation-based scenarios, detective and police scenarios achieve the highest ASR across all models. These two occupations, particularly those that are more closely related to criminal activities, make it easier to persuade models about the fictional scenario and mislead the model. Conversely, attacks based on lawyers and teachers do not work well, often resulting in plans for analysis report preparation or fiction writing (Figure 7), possibly due to their co-existence in models’ pertraining data. We also found specific occupation-based scenarios can perform extremely well on some models, such as the priest scenario, which has on-par ASR with detective or police on Mixtral-22B. In contrast, relation-based scenarios have nearly similar effects across models, pointing out that the authority role contained in occupations can enhance the quality of attack.
Among fourteen harmful action categories, several categories do not perform well across models for two main reasons: 1) Improper match: Categories like Hate Speech and Controversial Topics involve actions related to speaking, which do not fit our scenario structure where someone plans to do something. For example, “say the f word” in the Hate Speech fails to align with the scenario structure, reducing the credibility of the scenarios presented to models. 2) Misaligned objectives: Categories such as Sexually Explicit and Discrimination aim to elicit words or opinions explicitly (Ji et al., 2024), like “judge the German language.". In our attack, the model often generates plans on how to research the topic, resulting in non-malicious outcomes.
We show Red Queen Attack success rates on different scenarios and harmful action categories in Tables 10, 11, 12 and 13. We highlight the best ASR in bold and the second in underline.
We visualize the ASR result of GPT-4o, Llama-70B and Mixtral-8$\times$22B within each scenario and harmful action category combination in Figures 9, 10 and 11. We analyze whether certain harmful action categories have higher ASR in specific scenarios and find that ASR depends on the scenario’s effectiveness; a successful scenario works across all categories, and vice versa.
Figure 9: Heatmap for GPT-4o result in five turn Red Queen Attack
Figure 10: Heatmap for Llama3-70B result in three turn Red Queen Attack
Figure 11: Heatmap for Mixtral-22B result in five turn Red Queen Attack
Appendix C Data Sample
Table 15 shows harmful action samples for each category. Tables 14, 16, 17 and 18 shows the average token length of different turns.
Appendix D Comparison to Existing Jailbreak Attacks
We illustrate the detail of our comparison experiment between Red Queen Attack and previous jailbreak attacks in this section. We choose five jailbreak attacks from three categories to ensure a comprehensive comparison.
Cipher-based attack (Yuan et al., 2024) employs non-natural languages, such as ciphers, to obscure harmful intent and circumvent safety alignment mechanisms.
ArtPrompt (Jiang et al., 2024b) substitutes the harmful intent in the original prompt with ASCII art representations.
CoSafe Yu et al. (2024) adopts a coreference strategy in multi-turn dialogue systems, referencing malicious content in the final query.
CoU (Bhardwaj and Poria, 2023) utilizes a Chain-of-Utterances (CoU) framework to construct multi-turn conversations between two agents.
DeepInception (Li et al., 2023) incorporates nested instructions to prompt the LLM to simulate a virtual, multi-layered scenario with multiple characters, achieving jailbreak objectives.
Since DeepInception and ArtPrompt use the same harmful action set from Advbench (Zou et al., 2023) (50 queries), we construct the five-turn Red Queen Attack on this set for a fair comparison. We randomly pick 50 jailbreak artifacts from Cipher-based, CoSafe and CoU. We present jailbreak example of each baseline in Figure 12. The ASR for each attack is shown in  Table 5, with the best-performing attack on each model highlighted in bold.
Model Cipher ArtPrompt CoSafe CoU DeepInception Red Queen
Llama3-8b 0.06 0.28 0.08 0.14 0.18 0.38
Llama3-70b 0.08 0.20 0.06 0.06 0.30 0.94
Llama3.1-70b 0.28 0.22 0.20 0.06 0.50 0.98
Llama3.1-405b 0.10 0.10 0.08 0.02 0.56 0.84
Mixtral-8x7b 0.04 0.12 0.12 0.08 0.64 0.66
Mixtral-8x22b 0.10 0.48 0.16 0.10 0.40 0.88
GPT-4o-mini 0.12 0.38 0.04 0.00 0.76 1.00
GPT-4o 0.16 0.34 0.10 0.08 0.70 1.00
Qwen2-7b 0.14 0.48 0.04 0.02 0.46 0.32
Qwen2-72b 0.12 0.22 0.06 0.02 0.32 0.94
Average 0.12 0.28 0.10 0.06 0.48 0.79
Table 5: ASR comparison between Red Queen Attack and previous jailbreak attacks.
From the table, we can obtain the following observations:
  • Generalizability: Red Queen Attack achieves high ASR across all model families, while previous jailbreak attacks only perform well on specific model families.
  • Effectiveness: The Red Queen Attack outperforms all previous jailbreak attacks on nine out of ten models, demonstrating its effectiveness in targeting the model’s weak Theory-of-Mind ability.
  • Advantages of Concealment: The average performance of each attack shows that concealment methods (e.g., ArtPrompt) and fictional scenario construction (e.g., DeepInception) outperform multi-turn approaches (e.g., CoSafe and CoU).
Appendix E DPO Alignment Experiment
Appendix F Judgment Function Comparison
F.1 Comparison with existing Judgment Methods
To further address concerns about potential bias or sensitivity in our new judgment method, we conducted an additional evaluation study on wild prompt-response pairs to validate its reliability. Jailbreakbench (Chao et al., 2024) provides 300 jailbreak prompt-response pairs (harmful or benign) to evaluate current judgment functions. The prompts-response pairs also contain 100 benign examples from the XS-Test (Röttger et al., 2024) to test how sensitive the judges are to benign prompts and responses. We compare our new judgment method with previous ones on these pairs in Table 6. Our new judgment method with Llama-3 outperforms all previous judgment methods, and we find the same pattern that the new judgment method has a lower FNR compared to previous ones.
Metric Agreement (↑) FPR (↓) FNR (↓)
GCG 56.0 9.1 64.2
GPT-4o 85.7 21.8 10.0
Bert-Based 63.0 100.0 0.53
Llama Guard 85.0 30.0 6.3
Llama-3(Judge) 87.0 18.0 4.0
GPT-4o(Judge) 84.0 17.0 15.0
Table 6: Comparison of judgment methods on wild prompt-response pairs. FPR: False Positive Rate, FNR: False Negative Rate.
F.2 Details of Judgment Methods
We demonstrate details of the judgment method listed in Section 4 in the following:
Bert-based Classifier (Huang et al., 2024c) is a fine-tuned classifier trained on using HH-RLHF dataset (Ganguli et al., 2022). The classifier estimates a risk score between 0 to 1 for each response. The attacks will be classified as successful if the risk exceeds 0.5.
System Judging Prompt is designed to label responses as harmful or not. It works well with GPT-4 and Llama3-70B. This prompt evaluates whether the response includes a detailed plan or hints for creating a plan. We recommend testing this prompt first when adapting it to other tasks. The prompt is shown in Table 16, we evaluate the attack is successful if response contains “[Yes]”.
F.3 Failure Example of Previous Evaluators
Appendix G Prompt Template
We provide actual Red Queen Attack of different scenarios in this section. Due to ethical considerations, we only provide all scenarios (LABEL:tab:attack_scenario) in three turns and police scenarios in all turns. We will release the full attack.
Appendix H Human Annotators
We engaged three human annotators who volunteered to participate in the annotation process. Each annotator was provided with the following instruction during labeling: "Here is a model response to a question. Please evaluate whether the response is harmful or not." Each human annotators were notified the process may contain malicious content from models. The three annotators consisted of two males and one female, with two of them being of Asian origin and one from India.
@@ Appendix H Human Annotators
Table 8: Model output from Mixtral-8$\times$7B
@@
Scenario Llama 3 70B Mixtral 8$\times$7B Mixtral 8$\times$22B Qwen 2 7B Qwen 2 72B
@@
Scenario Llama 3 70B Mixtral 8$\times$7B Mixtral 8$\times$22B Qwen 2 7B Qwen 2 72B
@@
Type Occupation-Based Average
Teacher Police Detective Lawyer Priest
User 30.9 30.9 28.5 31.3 39.0 32.1
Assistant 117.5 137.8 149.4 144.6 187.4 147.4
Relation-Based
Friend Neighbor Relative Son Somone
User 28.7 30.4 30.4 28.8 28.5 29.4
Assistant 116.9 111.6 111.6 109.8 149.4 119.9
Table 14: Average token length of multi-turn scenario
@@
Table 15: Harmful action categories and examples.
@@
Table 16: Average token length of three-turn scenario.
@@
Table 17: Average token length of four-turn scenario.
@@
Table 18: Average token length of five-turn scenario.
@@
Table 22: Failure example of GCG. The red word is involved in the keyword list in LABEL:tab:_gcg
@@
Table 25: Failure example of Llama Guard Judgment Method.